# A Novel Skip Orthogonal List for Dynamic Optimal Transport Problem

### 05 Jan 2024 v4

@@ 1.1 Related Works
Approximate Algorithms. For approximate optimal transport, sherman2017generalized [25] proposed a $(1+\epsilon)$ approximation algorithm in $\epsilon^{-2}|E|^{1+o(1)}$ time.  pele2009fast [22] introduced the FastEMD algorithm that applies classic algorithms on a heuristic sketch of the input graph. Later, cuturi2013sinkhorn [9] used Sinkhorn-Knopp iterations to approximate the optimal transport problem by adding the smoothed entropic entry as the regularization term. Recently several optimizations on the Sinkhorn algorithm have been proposed, such as the Overrelaxation of Sinkhorn  [31] and the Screening Sinkhorn algorithm  [1].
Dynamic Trees. Our proposed data structure incorporates the techniques from dynamic trees, which can maintain and manipulate tree-like structures under specific operations. sleator1981data [26] proposed the Link/Cut tree data structure for path aggregate maintenance. tarjan1984finding [30] introduced the Euler Tour Technique for maintaining a tree as a cyclic ordered set of its Euler Tour sequence.
@@ 2.2 Euler Tour Technique
Through Definition 1, we can reduce edge linking, edge cutting, sub-tree weight modification and sub-tree weight querying to constant number of element insertion, element deletion, range weight modification and range weight querying on a circular ordered set [29]. We show in Section 4.1 that the dynamic optimal transport can be reduced to the 2D version of these four operations.
@@ 3 Skip Orthogonal List
For the remaining of the section, we construct a data structure with the expected $O(\absolutevalue{V}^{2})$ space complexity, where each operation can be done with the expected $O(\absolutevalue{V})$ time. Section 3.1 shows the overall structure of the data structure and how to query in this data structure. Section 3.2 illustrates the cut operation as an example based on this structure. For other operations (linking, insertion, and range updating), we leave them to appendix B.
@@ 3.1 The Overall Structure
Note that for any pair $u,v\in S$, we use $(u,v)$ to denote the corresponding element in $S\times S$; with a slight abuse of notations, we also use $(u,v)$ to denote the corresponding node at the $\ell$-th level in the Skip Orthogonal List. If $\ell$ is not specified, $(u,v)$ refers to the node at the bottom level.
@@
For any positive integer $n$, in a Skip Orthogonal List $\mathcal{L}$ over the cyclic ordered set $S$, suppose $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are 2 elements in $S\times S$. We say the node $(u_{1},v_{1})$ dominates $(u_{2},v_{2})$ at level $l$ if and only if the following three conditions are all satisfied:
  • $h(u_{1})\geq l$ and $h(v_{1})\geq l$;
  • $u_{1}=u_{2}$ or $\max_{u=u_{1}+1}^{u_{2}}h(u)<l$;
  • $v_{1}=v_{2}$ or $\max_{v=v_{1}+1}^{v_{2}}h(v)<l$.
@@ 3.2 The Update Operation: Cut
The generalized lazy propagation to our 2D data structure ensures that only the nodes that are “close” to the two rows and columns are modified, and consequently the updating time is guaranteed to be low. Specifically, the expected time complexity is $O(\absolutevalue{V})$ upon each call of the cut subroutine.
@@ 4 Our Dynamic Network Simplex Method
The simplex method performs simplex iterations on some initial feasible basis until the optimal solution is obtained. The simplex iterations are used for refining the current solution under the dynamic changes. In each simplex iteration, some variable with negative simplex multiplier is selected to perform a pivot subroutine, where one common strategy is to pivot in the variable with the smallest simplex multiplier. In Section 4.1 we focus on defining the dynamic optimal transport operations and using simplex iterations to solve this problem, while in Section 4.2 we analyze the details in each simplex iteration. Our method is presented in the context of the conventional Network Simplex algorithm [8, 21].
@@ 4.1 Dynamic Optimal Transport Operations
In an Optimal Transport problem, suppose the nodes in node set $V$ are located in some metric space $\mathcal{X}$, e.g., the Euclidean Space $\mathbb{R}^{k}$. The edge cost $c_{uv}$ is usually defined as the (squared) distance between $u$ and $v$ in the space. Let $w:\mathbb{R}^{V}$ denote the weight vector as defined in the equation (4). A Dynamic Optimal Transport algorithm should support the following four types of update as well as online query:
@@ 4.2 The Details for Simplex Iteration
    Update Primal Solution. Adding the new variable to the current basis forms a cycle. We send the circular flow in the cycle, until some basic variable in the reverse direction runs out of flow. Denote that node as $x_{i_{\rm out}j_{\rm out}}$, which is to be pivoted out.
@@
Input: Adjusted cost matrix $C^{\pi}$ corresponding to old solution $\pi$, entering variable $x_{i_{\rm in}j_{\rm in}}$, leaving variable $x_{i_{\rm out}j_{\rm out}}$.
Output: Updated adjusted cost matrix $C^{\pi^{\prime}}$ corresponding to new dual solution $\pi^{\prime}$.
1:  $t\leftarrow c^{\pi}_{i_{\rm in}j_{\rm in}}$.
2:  Cut the edge $\quantity(i_{\rm out},j_{\rm out})$ in $C^{\pi}$ and denote the 4 pieces as $C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}$.
3:  Range update on $C^{\pi}_{12}$ with increasing all the entries by $-t$.
4:  Range update on $C^{\pi}_{21}$ with increasing all the entries by $t$.
5:  Link the pieces $\{C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}\}$ by the edge $x_{i_{\rm in}j_{\rm in}}$, and obtain the adjusted $C^{\pi^{\prime}}$ as the output.
@@ 5 Experiments
Datasets. We study the performance of our algorithm on both synthetic and real datasets. For synthetic datasets, we construct a mixture of two Gaussian distributions in $\mathbb{R}^{2}$, where the points of the same Gaussian distribution share the same label. We also use the popular real-world dataset MNIST [18]. We partition the labels into two groups and compute the optimal transport between them.
Setup. We set the Sinkhorn regularization parameter $\lambda$ as $0.1$ and scale the median of the cost matrix $C$ to be 1. We vary the the node size |$\absolutevalue{V}$ | up to $4\times 10^{4}$. For each dataset, we test the static running time of POT on our machine and executed each dynamic operations 100 times to calculate the means and standard errors of our algorithm. For Spatial Position Modification, we randomly choose some point and add a random Gaussian noise with a variance of 0.5 in each dimension to it. For Point Insertion, we randomly select a point in the dataset that is outside the current OT instance to insert. We perform a query after these updates and compare with the static algorithms implemented in POT.
@@ 6 Conclusion and Future Work
Appendix A Space Complexity of Skip Orthogonal List
Theorem A.1
@@ Appendix A Space Complexity of Skip Orthogonal List
By Lemma 1, 2 and union bound, as $p\in(0,1)$ is a constant, we can see that there exists some constant $c_{1},c_{2}$ such that $N<c_{1}n^{2}$ with probability at least $1-\frac{c_{2}}{n}$, which finishes the proof of Theorem A.1.
Appendix B Detailed Update Algorithms on Skip Orthogonal List
This section provides detailed algorithm for our Skip Orthogonal List. We use the notation in Figure 4(a) and denote $a^{(i)}$ as the $i^{\rm th}$ ancestor of node $a$.
B.1 Cut
@@
0:  Skip 2D Orthogonal List $C^{\pi}$, Edge $\quantity{u,v}$ to cut
0:  Shattered $C^{\pi}$ after cutting the edge $\quantity{u,v}$
1:  Function $\text{cut}(C^{\pi},u,v)$
2:  {Find the 4 intersections of row $(u,v)$, row $(v,u)$ and column $(u,v)$, column $(v,u)$. Denote them as $p_{11},p_{12},p_{21},p_{22}$ as Figure 4(a) demonstrates.}
3:  Run 4 copies of push_down_line subroutine to clear the tag attribute of relevant nodes.
4:  Run 4 copies of cut_line subroutine to break the entire 2D Orthogonal List $C^{\pi}$ into 4 pieces.
5:  Run 4 copies of pull_up_lines subroutines to update the min attributes of affected nodes.
6:  return $C^{\pi}$
7:  EndFunction
[…]
9:  for Each $x$ in the Euler Tour of the current sub-tree do
10:     Move $p,q$ in the current direction.
11:     Let $p_{b},q_{b}$ be the backward node in the orthogonal direction and $p_{f},q_{f}$ be the forward node.
12:     Iterative calculate $p_{b}^{(i)},p_{f}^{(i)},q_{b}^{(i)},q_{f}^{(i)}$, i.e., the $i$-th ancestor.
13:     for $i=h_{x},h_{x}-1,\cdots,0$ do
14:        Link $p_{b}^{(i)}$ and $q_{f}^{(i)}$. Update parent attributes.
15:        Link $q_{b}^{(i)}$ and $p_{f}^{(i)}$. Update parent attributes.
16:        Call push_down subroutine to update the min and tag attribute of their children.
17:     end for
@@
19:  EndFunction
20:  Function pull_up_lines$(C^{\pi},p)$
21:  for Each Direction dir do
22:     for Each edge $x$ in the Euler Tour of the current direction do
23:        for $i=1,2,\cdots,h_{x}$ do
24:           Let $p^{(i)}$ be the $i^{\rm th}$ ancestor of $p$.
25:           Call pull_up subroutine to update the min attribute of $p^{(i)}$ by brutally traversing through its children set.
26:        end for
27:        Let $p$ be the next node in the Euler Tour.
28:     end for
@@
    Note that when linking adjacent nodes, we need to update the parent attributes of nodes $q_{f}^{(i)}$ and $p_{f}^{(i)}$ accordingly, and nodes below it accordingly. However, every node is affected at most once during the entire cut_line process, and the number of nodes affected are sure to be affected by pull_up_lines process.
  • Subroutine push_down_lines is capable of pushing down the tag attribute of all nodes alongside the rows and columns of the four nodes, i.e. the red nodes in 4(b).
  • Similar to push_down, subroutine pull_up_lines updates the min attribute of the red nodes in 4(b).
  • In subroutine push_down, $q$ is in the children set of $p$ if and only if $p$ is the parent of $q$. Later we will use this concept to analyze the time complexity.
  • Subroutine push_down_2_lines is similar to subroutine pull_up_line. Instead of calling push_down subroutine which updates parents’ min attribute from lower level to higher level, pull_up_line calls the pull_up subroutine which clears the tag attribute from higher level to lower level. It pushes down the row and column of the input node and the row and column next to the given node.
From analyses above, we can now show Theorem B.1.
Theorem B.1
The expected time complexity of subroutine cut is $O(\absolutevalue{V})$.
@@
From Algorithm 2, each push_down_lines, cut_line, pull_up_lines subroutine only apply constant number of modifications and queries on children of nodes alongside the 2 cut rows and columns , i.e. the red and transparent nodes and their children in Figure 4(b).
We first calculate the expected number of nodes removed in subroutine cut, i.e. the transparent nodes in Figure 4(b). The removed nodes in each level are in the row $\rho_{uv}$, $\rho_{vu}$ and column $\tau_{uv},\tau_{vu}$ where $\rho_{uv}\triangleq\quantity{\bigl{(}(u,v),e\bigr{)}:e\in E(D_{T})}$ and $\tau_{uv}\triangleq\quantity{\bigl{(}e,(u,v)\bigr{)}:e\in E(D_{T})}$. By the definition of node height, node $\bigl{(}(u,v),e\bigr{)}$, $\bigl{(}e,(u,v)\bigr{)}$, $\bigl{(}(v,u),e\bigr{)}$ and $\bigl{(}e,(v,u)\bigr{)}$ are in less than or equal to $h_{e}+1$ orthogonal lists. Therefore, the total number of removed points is less than or equal to $4\sum_{e\in E(D_{T})}(h_{e}+1)$. Thus the expected number of removed nodes is less than or equal to $4\sum_{e\in E(D_{T})}(\mathbb{E}(h_{e})+1)=\frac{4}{p}n=O(n)$.
Next we calculate the children of nodes whose the dominate set is affected after the cut update takes place, i.e. red nodes and their children after the cut update in Figure 4(b). Similar to counting transparent nodes, the expected number of red nodes is also less than or equal to $\frac{4}{p}n=O(n)$. Now we count the number of their children.
Recall that in Figure 4(a) we denote the 2 diagonal pieces of the big skip list as $C^{\pi}_{11}$ and $C^{\pi}_{22}$ while we denote the 2 non-diagonal pieces of the big skip list as $C^{\pi}_{21}$ and $C^{\pi}_{12}$. We count the number of red nodes and their children in the 4 pieces after the cut operation respectively. Denote $g(n)$ as the number of nodes in $C^{\pi}_{11}$ whose parent’s children set is affected, i.e. whose parent is a red node as Figure 4(b) demonstrates, when $n=\absolutevalue{E(D_{T_{1}})}$ |. Then $g(\absolutevalue{E(D_{T_{2}})})$ is that of $C^{\pi}_{22}$, and we know that:
@@
For $k=0,1,\cdots,n$, with probability $p^{k}(1-p)^{n-k}\binom{n}{k}$, the upper level contains $k$ nodes, i.e. contain $g(k)$ nodes which is visited in the cut subroutine (denote $g(0)=0$). By the linearity of expectation, the following holds:
@@
We can prove by induction that $\mathbb{E}[g(n)]\leq\frac{4-2p}{(1-p)^{2}}n$. Therefore, the expected number of visited nodes in the cut process (i.e. red nodes and their children in Figure 4(b)) in the 2 diagonal pieces $C^{\pi}_{11}$ and $c^{\pi}_{12}$ does not exceed $\mathbb{E}[g(\absolutevalue{E(D_{T_{1}})})]+\mathbb{E}[g(\absolutevalue{E(D_{T_{2}})})]\leq\frac{4-2p}{(1-p)^{2}}\absolutevalue{E(D_{T})}=O(\absolutevalue{E(D_{T})})$.
@@
B.2 Link
The link subroutine is very similar to the cut subroutine. It can be described as Algorithm 3.

### 27 Jan 2024 v5

@@ 1.1 Related Works
Approximate Algorithms. For approximate optimal transport, Sherman [25] proposed a $(1+\epsilon)$ approximation algorithm in $\epsilon^{-2}|E|^{1+o(1)}$ time. Pele and Werman [22] introduced the FastEMD algorithm that applies classic algorithms on a heuristic sketch of the input graph. Later, Cuturi [9] used Sinkhorn-Knopp iterations to approximate the optimal transport problem by adding the smoothed entropic entry as the regularization term. Recently several optimizations on the Sinkhorn algorithm have been proposed, such as the Overrelaxation of Sinkhorn  [30] and the Screening Sinkhorn algorithm  [1].
@@ 2.2 Euler Tour Technique
Through Definition 1, we can reduce edge linking, edge cutting, sub-tree weight updating and sub-tree weight querying to constant number of element insertion, element deletion, range weight modification and range weight querying on a circular ordered set [29]. We show in Section 4.1 that the dynamic optimal transport can be reduced to the 2D version of these four operations.
@@ 3 Skip Orthogonal List
For the remaining of the section, we construct a data structure with the expected $O(\absolutevalue{V}^{2})$ space complexity, where each operation can be done with the expected $O(\absolutevalue{V})$ time. Section 3.1 shows the overall structure of the data structure and how to query in this data structure. Section 3.2 illustrates the cut operation as an example based on this structure. For other operations (linking, insertion, and range updating), we leave them to appendix H.
@@ 3.1 The Overall Structure
Note that for any pair $u,v\in S$, we use $(u,v)$ to denote the corresponding element in $S\times S$; with a slight abuse of notations, we also use “$(u,v)$ at level $\ell$” to denote the corresponding node at the $\ell$-th level in the Skip Orthogonal List. If level $\ell$ is not specified in the context, $(u,v)$ refers to the node at the bottom level.
@@
For any positive integer $n$, in a Skip Orthogonal List $\mathcal{L}$ over the cyclic ordered set $S$, suppose $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are 2 elements in $S\times S$. We say the node $(u_{1},v_{1})$ dominates $(u_{2},v_{2})$ at level $\ell$ if and only if the following three conditions are all satisfied:
  • $h(u_{1})\geq l$ and $h(v_{1})\geq\ell$;
  • $u_{1}=u_{2}$ or $\max_{u=u_{1}+1}^{u_{2}}h(u)<\ell$;
  • $v_{1}=v_{2}$ or $\max_{v=v_{1}+1}^{v_{2}}h(v)<\ell$.
@@ 3.2 The Update Operation: Cut
The generalized lazy propagation to our 2D data structure ensures that only the nodes that are “close” to the two rows and columns are modified, and consequently the updating time is guaranteed to be low. Specifically, the expected time complexity is $O(\absolutevalue{V})$ upon each copy of procedure Cut.
@@ 4 Our Dynamic Network Simplex Method
The simplex method performs simplex iterations on some initial feasible basis until the optimal solution is obtained. The simplex iterations are used for refining the current solution under the dynamic changes. In each simplex iteration, some variable with negative simplex multiplier is selected for a copy of procedure Pivot, where one common strategy is to pivot in the variable with the smallest simplex multiplier. In Section 4.1 we focus on defining the dynamic optimal transport operations and using simplex iterations to solve this problem, while in Section 4.2 we analyze the details in each simplex iteration. Our method is presented in the context of the conventional Network Simplex algorithm [8, 21].
@@ 4.1 Dynamic Optimal Transport Operations
In an Optimal Transport problem, suppose the nodes in node set $V$ are located in some metric space $\mathcal{X}$, e.g., the Euclidean Space $\mathbb{R}^{k}$. The edge cost $c_{uv}$ is usually defined as the (squared) distance between $u$ and $v$ in the space. Let $w\in\mathbb{R}^{V}$ denote the weight vector as defined in the equation (4). A Dynamic Optimal Transport algorithm should support the following four types of update as well as online query:
@@ 4.2 The Details for Simplex Iteration
    Update Primal Solution. Adding the new variable to the current basis forms a cycle. We send the circular flow in the cycle, until some basic variable in the reverse direction runs out of flow, , through Graph Search (e.g. Depth First Search) or Link/Cut Tree [26]. Denote that node as $x_{i_{\rm out}j_{\rm out}}$, which is to be pivoted out.
@@
1:Adjusted cost matrix $C^{\pi}$ corresponding to old solution $\pi$, entering variable $x_{i_{\rm in}j_{\rm in}}$, leaving variable $x_{i_{\rm out}j_{\rm out}}$
2:Updated adjusted cost matrix $C^{\pi^{\prime}}$ corresponding to new dual solution $\pi^{\prime}$
3:$t\leftarrow c^{\pi}_{i_{\rm in}j_{\rm in}}$.
4:Cut the edge $\quantity(i_{\rm out},j_{\rm out})$ in $C^{\pi}$ and denote the 4 pieces as $C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}$.
5:Range update on $C^{\pi}_{12}$ with increasing all the entries by $-t$.
6:Range update on $C^{\pi}_{21}$ with increasing all the entries by $t$.
7:Link the pieces $\{C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}\}$ by the edge $x_{i_{\rm in}j_{\rm in}}$, and obtain the adjusted $C^{\pi^{\prime}}$ as the output.
@@ 5 Experiments
Datasets. We study the performance of our algorithm on both synthetic and real datasets. For synthetic datasets, we construct a mixture of two Gaussian distributions in $\mathbb{R}^{784}$, where the points of the same Gaussian distribution share the same label. We also use the popular real-world dataset MNIST [18]. We partition the labels into two groups and compute the optimal transport between them.
Setup. We set the Sinkhorn regularization parameter $\lambda$ as $0.1$ and scale the median of the cost matrix $C$ to be 1. We vary the the node size |$\absolutevalue{V}$ | up to $4\times 10^{4}$. For each dataset, we test the static running time of POT on our machine and executed each dynamic operations 100 times to calculate the means of our algorithm. For Spatial Position Modification, we randomly choose some point and add a random Gaussian noise with a variance of 0.5 in each dimension to it. For Point Insertion, we randomly select a point in the dataset that is outside the current OT instance to insert. We perform a query after these updates and compare with the static algorithms implemented in POT.
@@ 6 Conclusion and Future Work
G Space Complexity of Skip Orthogonal List
[…]
Theorem G.1
@@ G Space Complexity of Skip Orthogonal List
By Lemma 1, 2 and union bound, as $p\in(0,1)$ is a constant, we can see that there exists some constant $c_{1},c_{2}$ such that $N<c_{1}n^{2}$ with probability at least $1-\frac{c_{2}}{n}$, which finishes the proof of Theorem G.1.
H Operations on Skip Orthogonal List
This section provides detailed algorithm for operations on our Skip Orthogonal List. We use the notation in Figure 4(a) and denote $a^{(i)}$ as the $i^{\rm th}$ ancestor of node $a$.
H.1 Cut
@@
1:Skip 2D Orthogonal List $C^{\pi}$, Edge $\quantity{u,v}$ to cut
2:Shattered $C^{\pi}$ after cutting the edge $\quantity{u,v}$
3:procedure Cut($C^{\pi},u,v$)
4:     Find the 4 intersections of row $(u,v)$, row $(v,u)$ and column $(u,v)$, column $(v,u)$. Denote them as $p_{11},p_{12},p_{21},p_{22}$ as Figure 4(a) demonstrates.
5:     Run 4 copies of procedure PushDownLines to clear the tag attribute of relevant nodes.
6:     Run 4 copies of procedure CutLine to break the entire 2D Orthogonal List $C^{\pi}$ into 4 pieces.
7:     Run 4 copies of Procedure PullUpLines to update the min attributes of affected nodes.
8:end procedure
9:procedure CutLine($C^{\pi},p,q$)
10:     for each $x$ in the Euler Tour of the current sub-tree do
11:         Move $p,q$ in the current direction.
12:         Let $p_{b},q_{b}$ be the backward node in the orthogonal direction and $p_{f},q_{f}$ be the forward node.
13:         Iterative calculate $p_{b}^{(i)},p_{f}^{(i)},q_{b}^{(i)},q_{f}^{(i)}$, i.e., the $i$-th ancestor.
14:         for $i=h_{x},h_{x}-1,\cdots,0$ do
15:              Link $p_{b}^{(i)}$ and $q_{f}^{(i)}$. Update parent attributes.
16:              Link $q_{b}^{(i)}$ and $p_{f}^{(i)}$. Update parent attributes.
17:              Call procedure PushDown to update the min and tag attribute of their children.
@@
19:     end for
20:end procedure
21:procedure PullUpLines($C^{\pi},p$)
22:     for each Direction dir do
23:         for each edge $x$ in the Euler Tour of the current direction do
24:              for $i=1,2,\cdots,h_{x}$ do
25:                  Let $p^{(i)}$ be the $i^{\rm th}$ ancestor of $p$.
26:                  Call procedure PullUp to update the min attribute of $p^{(i)}$ by brutally traversing through its children set.
27:              end for
28:              Let $p$ be the next node in the Euler Tour.
@@
    Note that when linking adjacent nodes, we need to update the parent attributes of nodes $q_{f}^{(i)}$ and $p_{f}^{(i)}$ accordingly, and nodes below it accordingly. However, every node is affected at most once during the entire CutLine procedure, and the number of nodes affected are sure to be affected by PullUpLines process.
  • Procedure PushDownLines is capable of pushing down the tag attribute of all nodes alongside the rows and columns of the four nodes, i.e. the red nodes in 4(b).
  • Similar to PushDown, procedure PullUpLines updates the min attribute of the red nodes in 4(b).
  • In procedure PushDown, $q$ is in the children set of $p$ if and only if $p$ is the parent of $q$. Later we will use this concept to analyze the time complexity.
  • Procedure PushDownLines is similar to procedure PullUpLines. Instead of calling procedure PushDown which updates parents’ min attribute from lower level to higher level, PullUpLines calls procedure PullUp which clears the tag attribute from higher level to lower level. It pushes down the row and column of the input node and the row and column next to the given node.
From analyses above, we can now show Theorem H.1.
Theorem H.1
The expected time complexity of procedure Cut is $O(\absolutevalue{V})$.
@@
From Algorithm 2, each PushDownLines, CutLine, procedure PullUpLines only apply constant number of modifications and queries on children of nodes alongside the 2 cut rows and columns , i.e. the red and transparent nodes and their children in Figure 4(b).
We first calculate the expected number of nodes removed in procedure Cut, i.e. the transparent nodes in Figure 4(b). The removed nodes in each level are in the row $\rho_{uv}$, $\rho_{vu}$ and column $\tau_{uv},\tau_{vu}$ where $\rho_{uv}\triangleq\quantity{\bigl{(}(u,v),e\bigr{)}:e\in E(D_{T})}$ and $\tau_{uv}\triangleq\quantity{\bigl{(}e,(u,v)\bigr{)}:e\in E(D_{T})}$. By the definition of node height, node $\bigl{(}(u,v),e\bigr{)}$, $\bigl{(}e,(u,v)\bigr{)}$, $\bigl{(}(v,u),e\bigr{)}$ and $\bigl{(}e,(v,u)\bigr{)}$ are in less than or equal to $h_{e}+1$ orthogonal lists. Therefore, the total number of removed points is less than or equal to $4\sum_{e\in E(D_{T})}(h_{e}+1)$. Thus the expected number of removed nodes is less than or equal to $4\sum_{e\in E(D_{T})}(\mathbb{E}(h_{e})+1)=\frac{4}{p}n=O(n)$.
Next we calculate the children of nodes whose the dominate set is affected after procedure Cut update takes place, i.e. red nodes and their children after the Cut update in Figure 4(b). Similar to counting transparent nodes, the expected number of red nodes is also less than or equal to $\frac{4}{p}n=O(n)$. Now we count the number of their children.
Recall that in Figure 4(a) we denote the 2 diagonal pieces of the big skip list as $C^{\pi}_{11}$ and $C^{\pi}_{22}$ while we denote the 2 non-diagonal pieces of the big skip list as $C^{\pi}_{21}$ and $C^{\pi}_{12}$. We count the number of red nodes and their children in the 4 pieces after procedure Cut respectively. Denote $g(n)$ as the number of nodes in $C^{\pi}_{11}$ whose parent’s children set is affected, i.e. whose parent is a red node as Figure 4(b) demonstrates, when $n=\absolutevalue{E(D_{T_{1}})}$ |. Then $g(\absolutevalue{E(D_{T_{2}})})$ is that of $C^{\pi}_{22}$, and we know that:
@@
For $k=0,1,\cdots,n$, with probability $p^{k}(1-p)^{n-k}\binom{n}{k}$, the upper level contains $k$ nodes, i.e. contain $g(k)$ nodes which is visited in procedure Cut (denote $g(0)=0$). By the linearity of expectation, the following holds:
@@
We can prove by induction that $\mathbb{E}[g(n)]\leq\frac{4-2p}{(1-p)^{2}}n$. Therefore, the expected number of visited nodes in procedure Cut (i.e. red nodes and their children in Figure 4(b)) in the 2 diagonal pieces $C^{\pi}_{11}$ and $c^{\pi}_{12}$ does not exceed $\mathbb{E}[g(\absolutevalue{E(D_{T_{1}})})]+\mathbb{E}[g(\absolutevalue{E(D_{T_{2}})})]\leq\frac{4-2p}{(1-p)^{2}}\absolutevalue{E(D_{T})}=O(\absolutevalue{E(D_{T})})$.
@@
H.2 Link
Procedure Link is very similar to procedure Cut. It can be described as Algorithm 3.